# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

### v1

@@ Acknowledgements
This research was funded by Realta Fusion and performed as part of the DOE Milestone Based Fusion Development Program DE-SC0024887. The authors would like to thank the DOE Milestone review committee for their helpful input which improved this paper. This research also used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award FES-ERCAP0026655. Realta Fusion is a fusion energy company spun out of the ARPA-E funded WHAM project at the University of Wisconsin–Madison. Realta is developing compact magnetic mirror fusion power plants as the low capital cost path to fusion power, and targeting initial fusion power plant deployment at the $100-200~{}\mathrm{MW}$ scale at industrial sites and datacenters for process heat and/or electricity.
Appendix A The effect of electron temperature variation on end plug performance
In order to investigate the effect of increased $T_{e}$ and $\phi_{p}$ on ion confinement in simple mirror end plugs, modified versions of the CQL3D-m simulations which appeared in Section 4 were used. In these simulations, CQL3D-m was run with a Maxwellian electron background species with fixed temperature and an evolving electron density set equal to the ion density. A Pastukhov potential of $\phi_{p}=4T_{e}$ was set in all simulations. This is lower than the canonical $\phi_{p}\sim 5T_{e}$ result in simple mirrors (Pastukhov, 1974) and a result of the reduced $T_{i}/T_{e}$ ratio in the tandem mirror end plug which lowers the collisionality difference between the ions and electrons. The use of this somewhat reduced $\phi_{p}$ value is supported by the calculations of $\phi_{p}$ in the POPCONs analysis in Section 3 and the literature (Logan et al., 1980; Hua & Fowler, 2004).
To facilitate efficient scoping of the impact of electron temperature and Pastukhov potential variation on performance, a fixed end plug condition was used with $B_{0}=4~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $a_{m}=0.15~{}\mathrm{m}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$. The equilibrium in these cases was left static. This is roughly consistent with an assumption that the background value of $B_{0}$ would be adjusted such that the field after diamagnetism was accounted for was $4~{}\mathrm{T}$. The velocity-space resolution in these CQL3D-m simulations could be reduced to 300 points in $u_{0}$ and 200 points in $\vartheta$ as ions and electrons did not need to be coupled, and the Pastukhov potential calculation was turned off (these calculations need a refined grid as the thermal component of both distributions and the mirror loss boundaries must be finely resolved). Similarly, the realspace $\sqrt{\psi_{n}}$ flux surface resolution could be reduced to 14 as coupling to Pleiades was not performed, and as fewer flux surfaces were used in the simulation, the number of particles in the NBI source could also be reduced to $1\times 10^{5}$ (when flux surface volumes, particularly those of the inner surfaces, are larger it becomes much easier to consistently formulate a bounce-averaged NBI source).
Figure 10: Average density $\langle n\rangle$ versus the end plug electron temperature $T_{e}$ for an endplug with $B_{m}=25~{}\mathrm{T}$, $a_{m}=0.15~{}\mathrm{m}$, $R_{m}=6.25$, $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$, $\ell=4.5~{}\mathrm{m}$, $\theta_{\mathrm{NBI}}=45^{\circ}$, and $\phi_{p}=4T_{e}$.
Values of $T_{e}$ were scanned from $25-125~{}\mathrm{keV}$ in increments of $25~{}\mathrm{keV}$ and $\phi_{p}$ was varied in a corresponding fashion to examine the effect of operating at higher temperatures and potentials than those found in the simulations described in Section 4. In addition to varying $T_{e}$, the neutral beam energy was scanned from $240-600~{}\mathrm{keV}$ in increments of $120~{}\mathrm{keV}$. The results of this scan are shown in Figure 10 and show when higher values of $T_{e}$ are present from electron heating by alphas in the central cell of the tandem, higher energy neutral beams can be much more beneficial than they are in the simple mirror case examined in Section 4. The reason for this is two-fold: high energy ions are less susceptible to losses from being dragged or scattered into the ambipolar potential, and provided $E_{\mathrm{NBI}}$ is high enough for ions to not be promptly lost to the ambipolar potential, the end plug confinement improves at high $T_{e}$ to the extent that high energy beams build up large enough plasma densities to be absorbed efficiently. It should be noted here that these CQL3D-m results draw the electron drag dominated plug assumption made in (20) to simplify the tandem mirror POPCON system into question. Luckily however, the results of (20) should only vary weakly if the plug becomes more ion-ion scattering dominated. These results also indicate that the assumption that the increased end plug electron temperature in tandem mirrors should improve confinement if slightly higher energy neutral beams are used in Section 4 is accurate, and the general takeaways from the optimization in that section still hold; a viable end plug is achievable at modest beam powers and with realistic engineering assumptions.
Appendix B Limitations of RF ion heating in end plugs
@@
In previous high-field mirror literature, there has been some discussion of RF heating end plug ions to improve particle confinement (Forest et al., 2021; Endrizzi et al., 2023). However, it was decided to not use RF heating in this study. Over the course of this work, it became clear that an RF ion heating scheme like that employed in WHAM (Endrizzi et al., 2023), in which positive-ion, $\sim 100~{}\mathrm{keV}$, NBI ions are accelerated by $n=2-3$ harmonic ion-cyclotron heating, is not effective enough at increasing ion particle confinement, and at tandem mirror electron temperatures, positive-ion neutral beams with a beam energy of $120~{}\mathrm{keV}$ and an injection angle of $45^{\circ}$ will, according to (32), be nearly immediately lost to the ambipolar potential (this is supported by the analysis in Appendix A). In addition, particle losses from enhanced quasilinear diffusion into the loss-cone eliminate much of the gain in particle confinement that can be obtained from increasing the ion energy. While ion energy confinement and fusion rates are increased with the addition of RF heating, this is of negligible importance to overall tandem mirror fusion performance (and enhanced fusion rates in the end plug are in fact undesirable in most circumstances). In the remainder of this appendix, a representative RF heating case in at end plug conditions will be analyzed which demonstrate these challenges.
@@
The simulation under consideration here utilizes a tritium-fueled end plug with $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=5~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=120~{}\mathrm{keV}$, $P_{\mathrm{NBI}}=10~{}\mathrm{MW}$, $f_{\mathrm{RF}}=115.5~{}\mathrm{MHz}$, and $P_{\mathrm{RF}}=5~{}\mathrm{MW}$. This end plug condition is similar to the optimized end plugs identified in Section 4, but rather than using a $15~{}\mathrm{MW}$ of negative-ion NBI power, it uses $15~{}\mathrm{MW}$ combined of positive-ion NBI and $n=3$ harmonic RF power. The RF antenna in these simulations was placed slightly beyond the fast ion turning point at $z=1~{}\mathrm{m}$ calculated by CQL3D-m (this does not correspond exactly to $B/B_{0}=2$ due to the use of realistic neutral beam geometry and the effects of finite $\beta$) at the limiting flux surface $\sqrt{\psi_{n}}=1.0$, and the RF source frequency $f_{\mathrm{RF}}$ was set to correspond to a location just inside the fast ion turning point. The antenna spectrum used in GENRAY-c was consistent with the loop style of antennas which will be tested on WHAM (Endrizzi et al., 2023) and had a poloidal mode spectrum centered about $n_{\theta}=0$ with a spectral width of $n_{\theta}\pm 1$ and a uniform axial mode spectrum $n_{\phi}=0-10$ (this definition of the axial mode spectrum with $n_{\phi}$ is an artifact of the toroidal GENRAY code, but has been preserved in the text to make future replication with GENRAY-c easier). The simulations of this end plug were initially run with only NBI and no additional RF heating. CQL3D-m was iterated with Pleiades 8 times with 800 total CQL3D-m timesteps for $\sim 1~{}\mathrm{s}$ using only NBI heating until it reached steady-state. This was done in order to develop a target plasma with high enough $\beta$ to enable efficient cyclotron harmonic damping of the RF waves. When steady-state was achieved, GENRAY-c simulations were added to the CQL3D-m/Pleiades IPS workflow using self-consistent equilibrium and plasma profiles and coupled to CQL3D-m using an ion and electron quasilinear diffusion model. After the addition of RF, the simulations were run for an additional 1000 CQL3D-m time-steps, iterating Pleiades, CQL3D-m and GENRAY-c 8 times, for $\sim 1~{}\mathrm{s}$.
@@
The radial plasma $\langle E_{i}\rangle$ and $n$ profiles as well as the RF power deposition profiles from the simulations described above are shown in Figure 11. The addition of RF heating did provide a slight increase in the overall particle confinement and particle density, but the overall plug performance remained well below that which could be obtained using $15~{}\mathrm{MW}$ of negative-ion NBI (reference the plot shown in Figure 9 to see a comparable negative-ion NBI case). The majority of the improvement in particle confinement appears to have come from the $T_{e}$ increase due to the $\sim 1~{}\mathrm{MW}$ of electron directed RF heating by Landau damping. Ion directed heating’s effect on plasma performance was a mixed. There was a significant amount of inverse cyclotron harmonic damping that diffused particles downwards in velocity-space into the loss cone. Furthermore, quasilinear diffusion cannot increase the density anywhere in velocity-space to a larger value than it is at the neutral beam source location. Thus, particles at the neutral beam source energy continued to dominate the overall velocity-space pitch-angle scattering rate. Energy confinement increased, with the plasma reaching high average ion energies due to the formation of an RF driven energetic ion tail and quasilinear diffusion driven pumpout of lower energy ions, however, this is actually undesirable in an end plug. Just increasing average end plug particle energy without improving end plug particle confinement will not improve the tandem mirror central cell’s performance and will only serve to drive up $\beta$ putting the end plug closer to pressure driven instability limits. One potential solution would be to use higher harmonics which bias the magnitude of the quasilinear diffusion operator upwards in velocity space due to the quasilinear diffusion operator’s Bessel function dependence $Q\propto\mathrm{J}_{n}^{2}(k_{\perp}v_{\perp}/n\Omega_{i})$. However, harmonics greater than $n=3$ were found to be weakly damped in mirror end plug plasmas. Weak damping is undesirable in reactor scenarios as it can lead to poor antenna coupling and large in vessel fields that generate impurity sources. To increase damping rates, lower harmonics, $n=2,3$, had to be used. At lower harmonics, after the fast ion tail built up, damping became relatively strong, as shown in Figure 12, and the rays in the simulation damped after two passes through the resonance.
@@
As a result of simulations like the representative case shown here, RF ion heating was dropped from this study. The WHAM RF ion heating scheme was determined to not be scalable to tandem mirror fusion reactor end plugs. Positive-ion NBI cannot effectively fuel a tandem mirror end plug at relevant values of $T_{e}$ and $\phi_{p}$ and RF quasilinear diffusion at $n=2,3$ harmonics with this scheme was not effective enough at increasing particle confinement times. Finally, in cases of both weak and the strong damping there could potentially be validity issues with the raytracing approximation here. In cases of weak damping, interference and diffraction may impact the wave power deposition, and in cases of strong damping, the distance between the antenna and the wave resonance are comparable to the wavelength of the wave. Because of this, full-wave simulations may be needed to adequately model RF in magnetic mirrors due to the breakdown of the raytracing approximation. Future work may revisit RF ion heating in conjunction with negative ion neutral beams, modified antenna launch spectra which optimize the shape of quasilinear diffusion operator and the use of a full wave code such as AORSA (Jaeger et al., 2002).

### v2

@@ Acknowledgements
This research was funded by Realta Fusion and performed as part of the DOE Milestone Based Fusion Development Program DE-SC0024887. The authors would like to thank the DOE Milestone review committee for their helpful input which improved this paper. This research also used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award FES-ERCAP0026655. Realta Fusion is a fusion energy company spun out of the ARPA-E funded WHAM project at the University of Wisconsin–Madison. Realta is developing compact magnetic mirror fusion power plants as a low capital cost path to fusion power, and targeting initial fusion power plant deployment at the $100-200~{}\mathrm{MW}$ scale at industrial sites and datacenters for process heat and/or electricity.
Appendix A Limitations of RF ion heating in end plugs
@@
In previous high-field mirror literature, there has been some discussion of RF heating end plug ions to improve particle confinement (Forest et al., 2021; Endrizzi et al., 2023). However, it was decided to not use RF heating in this study. Over the course of this work, it became clear that an RF ion heating scheme like that employed in WHAM (Endrizzi et al., 2023), in which positive-ion, $\sim 100~{}\mathrm{keV}$, NBI ions are accelerated by $n=2-3$ harmonic ion-cyclotron heating, is not effective enough at increasing ion particle confinement, and at tandem mirror electron temperatures, positive-ion neutral beams with a beam energy of $120~{}\mathrm{keV}$ and an injection angle of $45^{\circ}$ will, according to (34), be nearly immediately lost to the ambipolar potential. In addition, particle losses from enhanced quasilinear diffusion into the loss-cone eliminate much of the gain in particle confinement that can be obtained from increasing the ion energy. While ion energy confinement and fusion rates are increased with the addition of RF heating, this is of negligible importance to overall tandem mirror fusion performance (and enhanced fusion rates in the end plug are in fact undesirable in most circumstances). In the remainder of this appendix, a representative RF heating case in at end plug conditions will be analysed which demonstrate these challenges.
@@
The simulation under consideration here utilizes a tritium-fueled end plug with $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=5~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=120~{}\mathrm{keV}$, $P_{\mathrm{NBI}}=10~{}\mathrm{MW}$, $f_{\mathrm{RF}}=115.5~{}\mathrm{MHz}$, and $P_{\mathrm{RF}}=5~{}\mathrm{MW}$. This end plug condition is similar to the optimized end plugs identified in Section 4, but rather than using a $15~{}\mathrm{MW}$ of negative-ion NBI power, it uses $15~{}\mathrm{MW}$ combined of positive-ion NBI and $n=3$ harmonic RF power. The RF antenna in these simulations was placed slightly beyond the fast ion turning point at $z=1~{}\mathrm{m}$ calculated by CQL3D-m (this does not correspond exactly to $B/B_{0}=2$ due to the use of realistic neutral beam geometry and the effects of finite $\beta$) at the limiting flux surface $\sqrt{\psi_{n}}=1.0$, and the RF source frequency $f_{\mathrm{RF}}$ was set to correspond to a location just inside the fast ion turning point. The antenna spectrum used in GENRAY-c was consistent with the loop style of antennas which will be tested on WHAM (Endrizzi et al., 2023) and had a poloidal mode spectrum centred about $n_{\theta}=0$ with a spectral width of $n_{\theta}\pm 1$ and a uniform axial mode spectrum $n_{\phi}=0-10$ (this definition of the axial mode spectrum with $n_{\phi}$ is an artifact of the toroidal GENRAY code, but has been preserved in the text to make future replication with GENRAY-c easier). The simulations of this end plug were initially run with only NBI and no additional RF heating. CQL3D-m was iterated with Pleiades 8 times with 800 total CQL3D-m timesteps for $\sim 1~{}\mathrm{s}$ using only NBI heating until it reached steady-state. This was done in order to develop a target plasma with high enough $\beta$ to enable efficient cyclotron harmonic damping of the RF waves. When steady-state was achieved, GENRAY-c simulations were added to the CQL3D-m/Pleiades IPS workflow using self-consistent equilibrium and plasma profiles and coupled to CQL3D-m using an ion and electron quasilinear diffusion model. After the addition of RF, the simulations were run for an additional 1000 CQL3D-m time-steps, iterating Pleiades, CQL3D-m and GENRAY-c 8 times, for $\sim 1~{}\mathrm{s}$.
@@
The radial plasma $\langle E_{i}\rangle$ and $n$ profiles as well as the RF power deposition profiles from the simulations described above are shown in Figure 12. The addition of RF heating did provide a slight increase in the overall particle confinement and particle density, but the overall plug performance remained well below that which could be obtained using $15~{}\mathrm{MW}$ of negative-ion NBI (reference the plot shown in Figure 9 to see a comparable negative-ion NBI case). The majority of the improvement in particle confinement appears to have come from the $T_{e}$ increase due to the $\sim 1~{}\mathrm{MW}$ of electron directed RF heating by Landau damping. Ion directed heating’s effect on plasma performance was a mixed. There was a significant amount of inverse cyclotron harmonic damping that diffused particles downwards in velocity-space into the loss cone. Furthermore, quasilinear diffusion cannot increase the density anywhere in velocity-space to a larger value than it is at the neutral beam source location. Thus, particles at the neutral beam source energy continued to dominate the overall velocity-space pitch-angle scattering rate. Energy confinement increased, with the plasma reaching high average ion energies due to the formation of an RF driven energetic ion tail and quasilinear diffusion driven pumpout of lower energy ions, however, this is actually undesirable in an end plug. Increasing average end plug particle energy without improving end plug particle confinement will not improve the tandem mirror central cell’s performance and will only serve to drive up $\beta$ putting the end plug closer to pressure driven instability limits. One potential solution would be to use higher harmonics which bias the magnitude of the quasilinear diffusion operator upwards in velocity space due to the quasilinear diffusion operator’s Bessel function dependence $Q\propto\mathrm{J}_{n}^{2}(k_{\perp}v_{\perp}/n\Omega_{i})$. However, harmonics greater than $n=3$ were found to be weakly damped in mirror end plug plasmas. Weak damping is undesirable in reactor scenarios as it can lead to poor antenna coupling and large in vessel fields that generate impurity sources. To increase damping rates, lower harmonics, $n=2,3$, had to be used. At lower harmonics, after the fast ion tail built up, damping became relatively strong, as shown in Figure 13, and the rays in the simulation damped after two passes through the resonance.
@@
As a result of simulations like the representative case shown here, RF ion heating was dropped from this study. The WHAM RF ion heating scheme was determined to not be scalable to tandem mirror fusion reactor end plugs, but will still be pursued experimentally on WHAM to quantify its effect on plasma performance and compared to simulations. Positive-ion NBI cannot effectively fuel a tandem mirror end plug at relevant values of $T_{e}$ and $\phi_{p}$ and RF quasilinear diffusion at $n=2,3$ harmonics with this scheme was not effective enough at increasing particle confinement times. Finally, in cases of both weak and the strong damping there could potentially be validity issues with the raytracing approximation here. In cases of weak damping, interference and diffraction may impact the wave power deposition, and in cases of strong damping, the distance between the antenna and the wave resonance are comparable to the wavelength of the wave. Because of this, full-wave simulations may be needed to adequately model RF in magnetic mirrors due to the breakdown of the raytracing approximation. Future work may revisit RF ion heating in conjunction with negative ion neutral beams, modified antenna launch spectra which optimize the shape of quasilinear diffusion operator and the use of a full wave code such as AORSA (Jaeger et al., 2002).